OpenStudy (anonymous):

Find the equation of a line with the given slope and y-intercept slope -5, y-intercept (0,-2)

OpenStudy (e.mccormick):

Know the slope intercept form of a line?

OpenStudy (anonymous):

I have tried and this is what I got y=-2/5x-2 do help me out

OpenStudy (e.mccormick):

\(y = mx + b\) Where \(m\) is the slope and \(b\) is the y-intercept.

OpenStudy (e.mccormick):

How did your slope change?

OpenStudy (anonymous):

I am new in math do help me solved it then i can learn it myself and do other problems

OpenStudy (e.mccormick):

You see the formula right there? \(y = mx + b\) \(m\) is slope \(b\) is intercept. In your problem they give you the slope and intercept. Replace the variables m and b with what you are given.

OpenStudy (anonymous):

Thats the formula I need the steps and solution

OpenStudy (e.mccormick):

That is the formula and the steps. All the steps are: put what was given in. Nothing else needs to be done for this one.

OpenStudy (anonymous):

Ask yourself this: What's the slope? What's the y-intercept? You already know the formula. correct? (y=mx+b) Since we know the that the slope is -5, and the y intercept is,(0,-2) or lets just say it's -2. to make it easier. Now fill those in the formula y=mx+b. What does m stand for? the slope What does b stand for? the y intercept. So just plug in -5 & -2 in the correct spots of the formula. If what I said is confusing please let me know. If not try plugging slope and y-intercept and I'll tell you if it's correct or not.